OpenStudy (babynini):

Using traces y=2x^2+z^2

OpenStudy (babynini):

x=0, y=z^2 parabola y=0, 0=2x^2+z^2 is..a circle? z=0, y=2x^2 which i thought is a parabola I think this one is all wrong heh

jimthompson5910 (jim_thompson5910):

The first and third statements are correct, but 0=2x^2+z^2 isn't a circle

OpenStudy (babynini):

I wasn't sure about that one. What is it? z=sqrt(2x^2) might be a better form..?

OpenStudy (babynini):

but turned in a different way since they have z and y^2 and I have z^2 and just y

jimthompson5910 (jim_thompson5910):

`but turned in a different way since they have z and y^2 and I have z^2 and just y` yes correct, swap y and z

OpenStudy (babynini):

hmhm excellent. I am still unsure of the steps to get the xz plane

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for 'z' \[\Large y = 2x^2+z^2\] \[\Large 0 = 2x^2+z^2\] \[\Large z^2 = -2x^2\] \[\Large z = \pm\sqrt{-2x^2}\] \[\Large z = \pm i*x \sqrt{2}\] So 'z' is not a real number for any nonzero x. However, if x were zero, then z would be zero as well. No other real number for x will make z a real number. So what this means is that the trace on the xz plane is simply a point. The point is at the origin (0,0). There is nothing else shown on this plane. Think of the very tip of the paraboloid touching the xz plane. I'm referring to something like this where the red point is touching the brown plane https://www.math.rutgers.edu/~greenfie/mill_courses/math251/gifstuff/min.gif

OpenStudy (babynini):

ahhh tricksy! Thanks so much for explaining that :D

jimthompson5910 (jim_thompson5910):

no problem